A 60 kg man is standing at the Earth's surface.

(a) What is the approximate magnitude of the gravitational force on the man by the Earth. Give your answer in newtons. Get the mass and radius of the Earth from the inside back cover of your text. Remember, enter only the number; don't try to enter the units. N

(b) What is the approximate magnitude of the force of gravitation on the same man by the Sun. You may approximate the distance from the Sun to the man by the distance from the Sun to the Earth; get that distance, as well as the mass of the Sun, from the inside back cover of your book. N

(c) What is the approximate magnitude of the force of gravitation on the same man by the Moon. You may approximate the distance from the Moon to the man by the distance from the Moon to the Earth; get that distance, as well as the mass of the Moon, from the inside back cover of your book. N
Answer #1

Ans:-

Given mass of man m2= 60kg

F = G m1m2/r^2

m1 = mass of earth = 6*10^24kg

r = distance between the center of earth and the man = radius of the earth

= 6.4*10^6m

G = 6.67*10^-11 Nm^2 /kg

F = (6.67*10^-11 * 6*10^24*60) /6.4*10^6

F= 586.2N

B]F = G m2*Ms /r^2

Rs= radius of sun = 1.496*10^11m

Re =6.38*10^6m

F = Gm2*Ms /(rs – Re)^2

= 6.67*10^-11*60*1.991*10^30 / (1.496*10^11 – 6.38*10^6)^2

F= 0.356N

C]F = G m2 Mm /(rm – re)^2

       = 6.67*10^-11*60*7.35*10^22/(384.400*10^6 – 6.38*10^6)^2

       = 0.002058N